# An Empirical Risk Minimization Approach for Offline Inverse RL and Dynamic Discrete Choice Model

### v4

@@ 7.2 Benchmark Algorithms
IQ-learn  IQ-learn is a popular gradient-based method, maximizing the occupancy matching objective (which does not guarantee that the Bellman equation is satisfied; see Web Appendix $S$B.6) for details.
BC  Behavioral Cloning (BC) simply minimizes the expected negative log-likelihood. This simple algorithm outperforms Zeng et al. (2023); Ziniu et al. (2022) many recent algorithms such as ValueDICE Kostrikov et al. (2019). See §$\S$4 for detailed discussions.
@@ 7.3.1 Performance results for the standard bus engine setting
We find that GLADIUS performs much better than non-Oracle baselines and performs at least on par with, or slightly better than, Oracle baselines. A natural question here is: why do the Oracle baselines that leverage the exact transition function and the precise linear parametrization not beat our approach? The main reason for this is the imbalance of state-action distribution from expert policy: (See Table 3 and Web Appendix §$\S$§A.1)
@@
    All trajectories start from mileage 1. In addition, the replacement action (action 0) resets the mileage to 1. Therefore, most states observed in the expert data are within mileage 1-5. When data collection policy visits a few states much more frequently than others, “the use of a projection operator onto a space of function approximators with respect to a distribution induced by the behavior policy can result in poor performance if that distribution does not sufficiently cover the state space.” (Tsitsiklis and Van Roy 1996a) This makes Oracle baseline predictions for states with mileage 1-5 slightly worse than GLADIUS.
@@ 8.1 Experimental Setup
For each baseline, we report the mean and standard deviation of 100 evaluation episodes after training. A higher % optimality indicates that the algorithm’s policy closely matches the expert. The 1000-episodic mean and standard deviation ([mean±$\pm$±std]) of the episodic reward of expert policy for each environment was $[232.77\pm 73.77]$ for Lunar-Lander v2 (larger the better), $[-82.80\pm 27.55]$ for Acrobot v1 (smaller the better), and $[500\pm 0]$ for Cartpole v1 (larger the better).
@@ A.1 More discussions on Bus engine replacement experiments
Figure A1, A2 and Table A1 - A4 shown below present the estimated results for reward and $Q^{\ast}$ using 50 trajectories (5,000 transitions) and 1,000 trajectories (100,000 transitions). As you can see in Figure A1 and A2, Rust and ML-IRL , which know the exact transition probabilities and employ correct parameterization (i.e., linear), demonstrate strong extrapolation capabilities for [Mileage, action] pairs that are rarely observed or entirely missing from the dataset (mileage 6-10). In contrast, GLADIUS, a neural network-based method, struggles with these underrepresented pairs. However, as we saw in the main text’s Table 2, GLADIUS achieves par or lower Mean Absolute Percentage Error (MAPE), which is defined as $\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\hat{r}_{i}-r_{i}}{r_{i}}\right|\times 100$ where $N$ is the total number of samples from expert policy $\pi^{*}$ and $\hat{r}_{i}$ is each algorithm’s estimator for the true reward $r_{i}$. This is because it overall outperforms predicting $r$ values for the [Mileage, action] pairs that appear most frequently and therefore contribute most significantly to the error calculation, as indicated by the visibility of the yellow shading in the tables below. (Higher visibility implies larger frequency.)
@@ B.1 Properties of Type 1 Extreme Value (T1EV) distribution
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ B.2 Properties of entropy regularization
When the regularizing function is constant $-\tau$ multiple of Shannon entropy $H(\mathbf{q})=-\sum_{i=1}^{|\mathcal{A}|}q_{i}\log\left(q_{i}\right)$,
@@
$\frac{\partial G}{\partial q_{i}\partial q_{j}}=\left\{\begin{array}[]{l}-\frac{\tau}{q_{i}},\quad\text{ if }i=j\\ 0,\quad\text{ otherwise. }\end{array}\right.$
This last equation states that the Hessian matrix is negative definite (since it is diagonal and $-\frac{\tau}{q_{1}}<0$ ), and thus ensures that the stationary point we compute is actually the maximum. Setting the gradient to $\mathbf{0}$ yields $q_{i}^{*}=\exp\left(\frac{x_{i}}{\tau}-1\right)$, however the resulting $\mathbf{q}^{*}$ might not be a probability distribution. To ensure $\sum_{i=1}^{n}q_{i}^{*}=1$, we add a normalization:
@@ B.5 Equivalence between DDC and Entropy regularized IRL
$Q$ that satisfies A10 is unique Rust (1994), and $Q-r$ forms a one-to-one relationship. Therefore, the exact equivalence between these two setups implies that the same reward function $r$ and discount factor $\beta$ will lead to the same choice-specific value function $Q$ and the same optimal policy for the two problems.
@@ B.6 IRL with occupancy matching
[…]
@@ C.1 Theory of TD correction using biconjugate trick
[…]
@@ C.2 Proof of Lemma 1
[…]
@@ C.3 Proof of Theorem 3
[…]
@@
[…]
=$=\left\{Q\in\mathcal{Q}\mid Q(s,a_{1})-Q(s,a_{2})=Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
@@
=$=\{Q\in\mathcal{Q}\mid D_{KL}(\pi^{*}(\cdot\mid s)\|\hat{p}_{Q}(\cdot\mid s))=0\text{ for all }s\in\bar{\mathcal{S}}\}$
=$=\{Q\in\mathcal{Q}\mid\hat{p}_{Q}(\cdot\mid s)=\pi^{*}(\cdot\mid s)\;\;\text{a.e.}\text{ for all }s\in\bar{\mathcal{S}}\}$
=$=\{Q\in\mathcal{Q}\mid\frac{\hat{p}_{Q}\left(a_{1}\mid s\right)}{\hat{p}_{Q}\left(a_{2}\mid s\right)}=\frac{\pi^{*}\left(a_{1}\mid s\right)}{\pi^{*}\left(a_{2}\mid s\right)}\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\}$
=$=\left\{Q\in\mathcal{Q}\mid\exp(Q(s,a_{1})-Q(s,a_{2}))=\exp\left(Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\right)\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
=$=\left\{Q\in\mathcal{Q}\mid Q(s,a_{1})-Q(s,a_{2})=Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
@@ C.4 Proof of Lemma 6
$\sigma_{\min}(M)=\sigma_{\min}\left(\tilde{M}\Sigma^{1/2}\right)\geq\sqrt{\lambda_{*}}\sigma_{\min}(\tilde{M})$
@@ C.5 Proof of Lemma 8 (Bellman error terms satisfying the PL condition)
$\;\Longleftarrow\log\left(\sum_{a^{\prime}\in\mathcal{A}}\exp Q_{\bm{\theta}}\left(s^{\prime},a^{\prime}\right)\right)\text{ is of }\mathcal{C}^{2}\quad(\because Q_{\bm{\theta}}\left(s^{\prime},a^{\prime}\right)\text{ is of }\mathcal{C}^{2}\text{ and its second derivative is bounded})$
$\;\Longleftarrow\;Q_{\bm{\theta}}\left(s,a\right)\text{ is of }\mathcal{C}^{2}\text{ for all }s\in\mathcal{S},a\in\mathcal{A}\quad(\because\text{logsumexp is of }\mathcal{C}^{\infty})$
@@
[…]
@@ C.6 Proof of Lemma 9 (NLL loss satisfying the PL condition)
$\frac{\partial f_{s,a}}{\partial Q_{\theta}(s,b)}=\pi_{\theta,b}-\mathbf{1}\{b=a\}:=g_{\theta,b},\quad\text{ and zero elsewhere. }$.
@@
$\;\Longleftarrow\;Q_{\bm{\theta}}\left(s,a\right)\text{ is of }\mathcal{C}^{2}\text{ for all }s\in\mathcal{S},a\in\mathcal{A}\quad(\because\text{logsumexp is of }\mathcal{C}^{\infty})$
@@
Define the averaged loss as ¯$\overline{\mathcal{L}}_{\!\text{NLL}}(\theta):=\mathbb{E}_{(s,a)\sim\mu}\bigl{[}f_{s,a}(\theta)\bigr{]}$, where $\mu$ is any probability measure on the product space $\mathcal{S}\times\mathcal{A}$.
Then, ¯$\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice continuously differentiable on $\Theta$, and its derivatives are given by
@@
The functions $\theta\mapsto f_{s,a}(\theta)$, $\theta\mapsto\nabla_{\theta}f_{s,a}(\theta)$, and $\theta\mapsto\nabla^{2}_{\theta}f_{s,a}(\theta)$ are continuous on $\Theta$ for each $(s,a)$. The norms of the first and second derivatives are dominated by the constants $M_{1}$ and $M_{2}$, respectively. Since constants are integrable with respect to any probability measure $\mu$, the conditions for the Leibniz integral rule are satisfied. Therefore, ¯$\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice differentiable and we can interchange differentiation and expectation:
@@
To establish that ¯$\overline{\mathcal{L}}_{\!\text{NLL}}$ is $\mathcal{C}^{2}$, we must show the Hessian is continuous. By the ‘Uniform $\mathcal{C}^{2}$ Bounds‘ assumption, the map $(\theta,s,a)\mapsto\nabla^{2}_{\theta}f_{s,a}(\theta)$ is continuous on $\Theta\times\mathcal{S}\times\mathcal{A}$. Let $(\theta_{n})_{n\in\mathbb{N}}$ be a sequence in $\Theta$ converging to $\theta\in\Theta$. Then for each $(s,a)$, $\nabla^{2}_{\theta}f_{s,a}(\theta_{n})\to\nabla^{2}_{\theta}f_{s,a}(\theta)$ as $n\to\infty$. The sequence of functions $g_{n}(s,a):=\nabla^{2}_{\theta}f_{s,a}(\theta_{n})$ is uniformly dominated in norm by the $\mu$-integrable constant $M_{2}$. By the Dominated Convergence Theorem, applied component-wise to the Hessian matrix,
@@
This confirms the continuity of the Hessian at any $\theta\in\Theta$. The same argument with the dominating function $M_{1}$ shows the gradient is continuous. Therefore, ¯$\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice continuously differentiable on $\Theta$. ∎

### v6

@@ 7.2 Benchmark Algorithms
IQ-learn  IQ-learn is a popular gradient-based method, maximizing the occupancy matching objective (which does not guarantee that the Bellman equation is satisfied; see Web Appendix SSB.6) for details.
BC  Behavioral Cloning (BC) simply minimizes the expected negative log-likelihood. This simple algorithm outperforms Zeng et al. (2023); Ziniu et al. (2022) many recent algorithms such as ValueDICE Kostrikov et al. (2019). See $\S$4 for detailed discussions.
@@ 7.3.1 Performance results for the standard bus engine setting
[…]
@@
    All trajectories start from mileage 1. In addition, the replacement action (action 0) resets the mileage to 1. Therefore, most states observed in the expert data are within mileage 1-5. When data collection policy visits a few states much more frequently than others, “the use of a projection operator onto a space of function approximators with respect to a distribution induced by the behavior policy can result in poor performance if that distribution does not sufficiently cover the state space.” (Tsitsiklis and Van Roy 1996b) This makes Oracle baseline predictions for states with mileage 1-5 slightly worse than GLADIUS.
@@ 8.1 Experimental Setup
[…]
@@ A.1 More discussions on Bus engine replacement experiments
Figure A1, A2 and Table A1 - A4 shown below present the estimated results for reward and $Q^{\ast}$ using 50 trajectories (5,000 transitions) and 1,000 trajectories (100,000 transitions). As you can see in Figure A1 and A2, Rust and ML-IRL , which know the exact transition probabilities and employ correct parameterization (i.e., linear), demonstrate strong extrapolation capabilities for [Mileage, action] pairs that are rarely observed or entirely missing from the dataset (mileage 6-10). In contrast, GLADIUS, a neural network-based method, struggles with these underrepresented pairs.
However, as we saw in the main text’s Table 2, GLADIUS achieves par or lower Mean Absolute Percentage Error (MAPE), which is defined as $\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\hat{r}_{i}-r_{i}}{r_{i}}\right|\times 100$ where $N$ is the total number of samples from expert policy $\pi^{*}$ and $\hat{r}_{i}$ is each algorithm’s estimator for the true reward $r_{i}$. This is because it overall outperforms predicting $r$ values for the [Mileage, action] pairs that appear most frequently and therefore contribute most significantly to the error calculation, as indicated by the visibility of the yellow shading in the tables below. (Higher visibility implies larger frequency.)
@@ B.1 Properties of Type 1 Extreme Value (T1EV) distribution
 $\begin{gathered}\text{ CDF: }F(x;\nu)=e^{-e^{-(x-\nu)}}\\ \text{ PDF: }f(x;\nu)=e^{-\left((x-\nu)+e^{-(x-\nu)}\right)}\\ \text{ Mean }=\nu+\gamma\end{gathered}$
@@
$P\left(G_{k}>\max_{i\neq k}G_{i}\right)$ $=\int_{-\infty}^{\infty}P\left(\max_{i\neq k}G_{i}<x\right)f_{G_{k}}(x)dx$
$=\int_{-\infty}^{\infty}e^{-\sum_{i\neq k}e^{-\left(x-\nu_{i}\right)}}e^{-\left(x-\nu_{k}\right)}e^{-e^{-\left(x-\nu_{k}\right)}}dx$
$=e^{\nu_{k}}\int_{-\infty}^{\infty}e^{-e^{-x}\sum_{i}e^{\nu_{i}}}e^{-x}dx$
@@
$=e^{\nu_{k}}\int_{0}^{\infty}e^{-uS}du=e^{\nu_{k}}\left[-\frac{1}{S}e^{-uS}\right]_{0}^{\infty}=\frac{e^{\nu_{k}}}{S}$
@@
Let $\nu_{1}-\nu_{2}=c$. Then $\mathbb{E}\left[G_{1}\mid G_{1}\geq G_{2}\right]$ is equivalent to $\nu_{1}+\frac{\int_{-\infty}^{+\infty}xF(x+c)f(x)\mathrm{d}x}{\int_{-\infty}^{+\infty}F(x+c)f(x)\mathrm{d}x}$, where the pdf $f$ and cdf $F$ are associated with $\text{Gumbel }(0,1)$, because
$P\left(G_{1}\leq x,G_{1}\geq G_{2}\right)$ $=\int_{-\infty}^{x}F_{G_{2}}(t)f_{G_{1}}(t)dt=\int_{-\infty}^{x}F\left(t-\nu_{2}\right)f\left(t-\nu_{1}\right)dt$
$\mathbb{E}\left[G_{1}\mid G_{1}\geq G_{2}\right]$ $=\frac{\int_{-\infty}^{\infty}xF\left(x+c-\nu_{1}\right)f\left(x-\nu_{1}\right)dx}{\int_{-\infty}^{\infty}F\left(x+c-\nu_{1}\right)f\left(x-\nu_{1}\right)dx}$
$=\frac{\int_{-\infty}^{\infty}\left(y+\nu_{1}\right)F(y+c)f(y)dy}{\int_{-\infty}^{\infty}F(y+c)f(y)dy}$
$=\nu_{1}+\frac{\int_{-\infty}^{\infty}yF(y+c)f(y)dy}{\int_{-\infty}^{\infty}F(y+c)f(y)dy}$
@@
$\begin{aligned} \int_{-\infty}^{+\infty}F(x+c)f(x)\mathrm{d}x&=\int_{-\infty}^{+\infty}\exp\{-\exp[-x-c]\}\exp\{-x\}\exp\{-\exp[-x]\}\mathrm{d}x\\ &\stackrel{{\scriptstyle a=e^{-c}}}{{=}}\int_{-\infty}^{+\infty}\exp\{-(1+a)\exp[-x]\}\exp\{-x\}\mathrm{d}x\\ &=\frac{1}{1+a}\left[\exp\left\{-(1+a)e^{-x}\right\}\right]_{-\infty}^{+\infty}\\ &=\frac{1}{1+a}\end{aligned}$
@@ B.2 Properties of entropy regularization
When the regularizing function is constant $\tau$ multiple of Shannon entropy $H(\mathbf{q})=-\sum_{i=1}^{|\mathcal{A}|}q_{i}\log\left(q_{i}\right)$,
@@
[…]
This last equation states that the Hessian matrix is negative definite (since it is diagonal and $\frac{\tau}{q_{1}}<0$ ), and thus ensures that the stationary point we compute is actually the maximum. Setting the gradient to $\mathbf{0}$ yields $q_{i}^{*}=\exp\left(\frac{x_{i}}{\tau}-1\right)$, however the resulting $\mathbf{q}^{*}$ might not be a probability distribution. To ensure $\sum_{i=1}^{n}q_{i}^{*}=1$, we add a normalization:
@@ B.5 Equivalence between DDC and Entropy regularized IRL
$Q$ that satisfies A10 is unique Rust (1994), and QrQ-r forms a one-to-one relationship. Therefore, the exact equivalence between these two setups implies that the same reward function $r$ and discount factor $\beta$ will lead to the same choice-specific value function $Q$ and the same optimal policy for the two problems.
@@ B.6 IRL with occupancy matching
$r(s_{0},a_{0})+\beta\cancel{\mathbb{E}_{s^{\prime}\sim P(s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[f(s^{\prime},a^{\prime})\mid s,a]}-f(s_{0},a_{0})$
$+\beta\mathbb{E}_{(s,a)\sim d_{1}^{\pi}}\left[r(s,a)+\beta\cancel{\mathbb{E}_{s^{\prime}\sim P(s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[f(s^{\prime},a^{\prime})\mid s,a]}-\cancel{f(s,a)}\mid s_{0},a_{0}\right]$
$+\beta^{2}\mathbb{E}_{(s,a)\sim d_{2}^{\pi}}\left[r(s,a)+\beta\cancel{\mathbb{E}_{s^{\prime}\sim P(s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[f(s^{\prime},a^{\prime})\mid s,a]}-\cancel{f(s,a)}\mid s_{0},a_{0}\right]$
@@ C.1 Theory of TD correction using biconjugate trick
$=2(\hat{\mathcal{T}}{Q}\rho-\hat{\mathcal{T}}{Q}Q-\cancel{Q\rho}+Q^{2})-(\rho^{2}-\cancel{2Q\rho}+Q^{2})$
$=2\hat{\mathcal{T}}{Q}\rho-2\hat{\mathcal{T}}{Q}Q+\cancel{2}Q^{2}-\rho^{2}-\cancel{Q^{2}}$
@@ C.2 Proof of Lemma 1
and define an operator $\mathcal{T}_{f}:\mathcal{F}\rightarrow$ $\mathcal{F}$ that acts on functions $f_{Q}$ :
@@ C.3 Proof of Theorem 3
$\left\{Q\in\mathcal{Q}\mid\hat{p}_{Q}(\;\cdot\mid s)=\pi^{*}(\;\cdot\mid s)\quad\forall s\in\bar{\mathcal{S}}\quad\text{a.e.}\right\}$
@@
$=\left\{Q\in\mathcal{Q}\mid\hat{p}_{Q}(\;\cdot\mid s)=\pi^{*}(\;\cdot\mid s)\quad\forall s\in\bar{\mathcal{S}}\quad\text{a.e.}\right\}$
$=\left\{Q\in\mathcal{Q}\mid Q(s,a_{1})-Q(s,a_{2})=Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
@@
$=\{Q\in\mathcal{Q}\mid D_{KL}(\pi^{*}(\cdot\mid s)\|\hat{p}_{Q}(\cdot\mid s))=0\text{ for all }s\in\bar{\mathcal{S}}\}$
$=\{Q\in\mathcal{Q}\mid\hat{p}_{Q}(\cdot\mid s)=\pi^{*}(\cdot\mid s)\;\;\text{a.e.}\text{ for all }s\in\bar{\mathcal{S}}\}$
$=\{Q\in\mathcal{Q}\mid\frac{\hat{p}_{Q}\left(a_{1}\mid s\right)}{\hat{p}_{Q}\left(a_{2}\mid s\right)}=\frac{\pi^{*}\left(a_{1}\mid s\right)}{\pi^{*}\left(a_{2}\mid s\right)}\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\}$
$=\left\{Q\in\mathcal{Q}\mid\exp(Q(s,a_{1})-Q(s,a_{2}))=\exp\left(Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\right)\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
$=\left\{Q\in\mathcal{Q}\mid Q(s,a_{1})-Q(s,a_{2})=Q^{*}(s,a_{1})-Q^{*}(s,a_{2})\quad\forall a_{1},a_{2}\in\mathcal{A},s\in\bar{\mathcal{S}}\right\}$
@@ C.4 Proof of Lemma 6
[…]
@@ C.5 Proof of Lemma 8 (Bellman error terms satisfying the PL condition)
[…]
@@
where the last inequality comes from assuming the dual norm is normalized such that $\left\|\delta_{(s,a)}\right\|_{H^{*}}=1$. [^16] Here, $H^{*}=(C(S\times A))^{*}$ is the space of Radon measures endowed with the total‐variation norm, and $\delta_{(s,a)}$ is the Dirac point‐mass. Substituting back, we get
@@ C.6 Proof of Lemma 9 (NLL loss satisfying the PL condition)
[…]
@@
[…]
@@
Define the averaged loss as $\overline{\mathcal{L}}_{\!\text{NLL}}(\theta):=\mathbb{E}_{(s,a)\sim\mu}\bigl[f_{s,a}(\theta)\bigr]$, where $\mu$ is any probability measure on the product space $\mathcal{S}\times\mathcal{A}$.
Then, $\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice continuously differentiable on $\Theta$, and its derivatives are given by
@@
The functions $\theta\mapsto f_{s,a}(\theta)$, $\theta\mapsto\nabla_{\theta}f_{s,a}(\theta)$, and $\theta\mapsto\nabla^{2}_{\theta}f_{s,a}(\theta)$ are continuous on $\Theta$ for each $(s,a)$. The norms of the first and second derivatives are dominated by the constants $M_{1}$ and $M_{2}$, respectively. Since constants are integrable with respect to any probability measure $\mu$, the conditions for the Leibniz integral rule are satisfied. Therefore, $\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice differentiable and we can interchange differentiation and expectation:
@@
To establish that $\overline{\mathcal{L}}_{\!\text{NLL}}$ is $\mathcal{C}^{2}$, we must show the Hessian is continuous. By the ‘Uniform $\mathcal{C}^{2}$ Bounds‘ assumption, the map $(\theta,s,a)\mapsto\nabla^{2}_{\theta}f_{s,a}(\theta)$ is continuous on $\Theta\times\mathcal{S}\times\mathcal{A}$. Let $(\theta_{n})_{n\in\mathbb{N}}$ be a sequence in $\Theta$ converging to $\theta\in\Theta$. Then for each $(s,a)$, $\nabla^{2}_{\theta}f_{s,a}(\theta_{n})\to\nabla^{2}_{\theta}f_{s,a}(\theta)$ as $n\to\infty$. The sequence of functions $g_{n}(s,a):=\nabla^{2}_{\theta}f_{s,a}(\theta_{n})$ is uniformly dominated in norm by the $\mu$-integrable constant $M_{2}$. By the Dominated Convergence Theorem, applied component-wise to the Hessian matrix,
@@
This confirms the continuity of the Hessian at any $\theta\in\Theta$. The same argument with the dominating function $M_{1}$ shows the gradient is continuous. Therefore, $\overline{\mathcal{L}}_{\!\text{NLL}}$ is twice continuously differentiable on $\Theta$. ∎